OpenStudy (anonymous):

A positive integer is one more than twice another. Their product is 17. find two integers.

OpenStudy (aravindg):

Let numbers be a,b a=1+2b ab=17

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what would be the next step?

OpenStudy (aravindg):

Substitute a from first equation. Then solve for b.

OpenStudy (aravindg):

(1+2b)b=17

OpenStudy (anonymous):

so 2b^2 +b-17=0 ???

OpenStudy (aravindg):

Yes find the roots. Only take the positive value among them as it is mentioned positive integer.

OpenStudy (anonymous):

great got it…thanks :)

OpenStudy (aravindg):

yw :)